1) A 60.0-kg student runs at a constant velocity up a flight of stairs. If the vertical distance of the stairs is 3.2 m, the work done against gravity is ___x10^3 J 


Use the information below to answer the following 4 questions.

A 10.0-kg object is accelerated horizontally from rest to a velocity of 11.0 m/s in 5.00 s by a horizontal force.


2) The acceleration of the object is ___m/s2


3) The horizontal force given to the object is ___N


4) The distance travelled by the object is ___m.


5) The work done on the object is ___J.




1
Expert's answer
2021-07-16T10:23:07-0400

"1)"

"\Delta h =3.2m"

"m = 60 kg"

"g = 9.8\frac{m}{s^2}"

"W = mg\Delta h = 60*9.8*3.2=1.88*10^3J"


"m_1= 10kg"

"v_0 = 0;v_1= 11\frac{m}{s}"

"t = 5s"

"2)"

"a = \frac{v_1-v_0}{t}= \frac{11-0}{5}=2.2\frac{m}{s^2}"

"3)"

"F = ma = 10*2.2 = 22 N"

"4)"

"s= v_0t +\frac{at^2}{2}= 0+\frac{2.2*25}{2}= 27.5m"

"5)"

"W_1 = Fs = 22*27.5 = 605J"

"\text{Answer:}"

"1)\text{...the work done against gravity is }1.88*10^3 J"

"\text{2) The acceleration of the object is }2.2m/s^2"

"\text{3) The horizontal force given to the object is }22N"

"\text{4) The distance travelled by the object is }27.5m"

"\text{5) The work done on the object is }605J"